# Negative cosmological constant in the dark energy sector: tests from JWST photometric and spectroscopic observations of high-redshift galaxies

### v1

@@ Abstract
Early observations with the James Webb Space Telescope (JWST) have revealed the existence of an unexpectedly large abundance of extremely massive galaxies at redshifts $z\gtrsim 5$: these observations are in tension with the predictions not only of the standard $\Lambda$CDM cosmology, but also with those of a wide class of dynamical dark energy (DE) models, and are generally in better agreement with models characterized by a phantom behaviour. Here we consider a model, inspired by string theory and the ubiquity of anti-de Sitter vacua therein, featuring an evolving DE component with positive energy density on top of a negative cosmological constant, argued in an earlier exploratory analysis to potentially be able to explain the JWST observations. We perform a robust comparison of this model against JWST data, considering both photometric observations from the CEERS program, and spectroscopic observations from the FRESCO survey. We show that the model is able to accommodate the JWST observations, with a consistency probability of up to $98\%$, even in the presence of an evolving component with a quintessence-like behaviour (easier to accommodate theoretically compared to phantom DE), while remaining consistent with standard low-redshift probes. Our results showcase the tremendous potential of measurements of high-redshift galaxy abundances in tests of fundamental physics, and their valuable complementarity with standard cosmological probes.
cosmology: cosmological models — cosmology: cosmological parameters — cosmology: cosmological parameters: dark energy — galaxies: abundances — galaxies: high-redshift galaxies
@@ 1 Introduction
The concordance $\Lambda$CDM cosmological model constitutes at present one of the simplest frameworks to describe the evolution of our Universe, and owes its success to its ability to describe a wide range of cosmological and astrophysical observations. One of the cornerstones of current cosmology is dark energy (DE, Peebles & Ratra, 2003; Frieman et al., 2008; Nojiri et al., 2017), a component with negative pressure and positive energy density driving the late-time acceleration of the Universe, whose discovery in 1998 led to one of the biggest paradigm shifts in physics (Riess et al., 1998; Perlmutter et al., 1999): within $\Lambda$CDM, DE takes the form of a positive cosmological constant (CC) $\Lambda$ (Carroll, 2001). Over the past two and a half decades, the nature of DE has to a large extent been investigated by means of three classes of cosmological observations: low-redshift ($z\lesssim 2$) probes based on Type Ia Supernovae (SNeIa) and Baryon Acoustic Oscillations (BAO), the latter observed in the clustering of various tracers of the large-scale structure, such as galaxies; and high-redshift ($z\sim 1100$) probes based on the Cosmic Microwave Background (CMB): see Huterer & Shafer (2018) for a recent review on observational probes of DE.  [^1] Nevertheless, it is worth emphasizing that the sensitivity of the CMB to (late-time) DE is indirectly mostly a low-redshift (geometrical) one, through the effect of DE on the distance to the CMB itself (see e.g. the recent discussion in Escamilla et al., 2023), whereas the sensitivity of the CMB to DE perturbations and the dynamics thereof is quite limited, except possibly in the case of models which deviate substantially from the positive CC $\Lambda$, which are however ruled out by observations. By virtue of the significant efforts devoted to the study of observations, underlying theory, and systematics associated to these probes, CMB, BAO, and SNeIa are generally referred to as standard cosmological probes and enjoy a (rightly deserved) privileged status for what concerns cosmological tests of the expansion of the Universe and of fundamental physics, with particular regard to the nature of DE. However, a series of growing cosmological tensions which appear to challenge the $\Lambda$CDM model (Perivolaropoulos & Skara, 2022), and the fact that some of these probes are close to being systematics-limited, are making it imperative to start looking for new independent probes, whose complementarity and synergy with the standard CMB+BAO+SNeIa can greatly enrich the landscape of methods to study the Universe: in this sense, a number of so-called “emerging” cosmological probes are rapidly gaining momentum, as beautifully summarized in the recent review of Moresco et al. (2022) [^2] Of course tests of DE are not limited to cosmological scales. Various tests of DE on astrophysical, astronomical, and even local scales, have been considered over the past decades, see Hamilton et al. (2015); Burrage & Sakstein (2016, 2018); Vagnozzi et al. (2021); Zhang (2022); Ferlito et al. (2022); Brax et al. (2023); Benisty & Capozziello (2023); Zhang & Zhang (2023); Benisty et al. (2023c, a, b); Paraskevas & Perivolaropoulos (2023); Benisty et al. (2023); Kaneta et al. (2023) for examples in this sense.
@@
In this context, the long-awaited recent observations of massive, distant galaxies delivered by the James Webb Space Telescope (JWST, Gardner et al., 2006) have opened a remarkable window not only onto astrophysics and the process of galaxy formation, but also onto novel cosmological tests of the contents of our Universe and more generally fundamental physics (Labbé et al., 2023; Xiao et al., 2023; Arrabal Haro et al., 2023). This is possible because the abundance of galactic haloes of dark matter (DM) mass $M$ at redshifts $z\approx 5-10$ is predicted to be exponentially sensitive to the growth factor of perturbations $D(z)$, while also being a strongly decreasing function of redshift, and one which depends crucially on the (model-dependent) time-redshift relation (Del Popolo & Yesilurt, 2007) [^3] Distant galaxies are also useful as a test of cosmological models through their ages, which of course cannot exceed the age of the Universe: see for instance Jimenez et al. (2019); Valcin et al. (2021); Bernal et al. (2021); Vagnozzi et al. (2022); Wei & Melia (2022); Cimatti & Moresco (2023); Costa et al. (2023) for recent works exploring the important of stellar and galactic ages as a test of cosmological models. On the other hand, if we denote by $\Omega_{b}$ and $\Omega_{m}$ the cosmic density parameters for baryons and the total matter component respectively (and with $f_{b}\equiv\Omega_{b}/\Omega_{m}$), the stellar mass of a galaxy $M_{\star}$ cannot exceed the maximal allowed baryonic content of the host halo $f_{b}M$. This implies that comparing the theoretically predicted cumulative comoving stellar number density of massive ($M_{\star}\approx 10^{10}-10^{11}\,M_{\odot}$) galaxies, $n(>M_{\star},z)$, with the observed abundance of galaxies of given stellar mass at redshifts $z\approx 5-10$, can potentially provide strong constraints on the expansion history and growth of cosmological perturbations, and therefore also on the nature of DE, in a way which is completely independent of the complex baryonic physics involved in the formation of galaxies. Needless to say, such constraints would be extremely valuable in light of their high complementarity to state-of-the-art determinations from CMB, BAO, and SNeIa, both in terms of range of redshifts/cosmic times probed, as well as potential systematics involved. As hinted to earlier, the possibility of testing the expansion history of the Universe at such intermediate redshifts is particularly valuable in view of the tension between local and high-redshift determinations of the Hubble constant $H_{0}$ (see e.g. Verde et al., 2019; Di Valentino et al., 2021b, c; Abdalla et al., 2022; Schöneberg et al., 2022; Shah et al., 2021; Kamionkowski & Riess, 2023; Hu & Wang, 2023; Vagnozzi, 2023; Verde et al., 2023, for recent reviews). At a significance of $\gtrsim 5\sigma$, the Hubble tension poses what is probably one of the most serious observational challenges to $\Lambda$CDM, and can potentially lead us to completely rethink the nature of DE.  [^4] With no claims as to completeness, see Yang et al. (2018a); Guo et al. (2019); Di Valentino et al. (2020b, c); Zumalacarregui (2020); Gómez-Valent et al. (2020); Di Valentino et al. (2020a); Gao et al. (2021); Petronikolou et al. (2022); Alestas et al. (2022); Roy et al. (2022); Sharma et al. (2022); Moshafi et al. (2022); Schiavone et al. (2023); Gao et al. (2022); Bernui et al. (2023); Ben-Dayan & Kumar (2023); Zhai et al. (2023); Basilakos et al. (2023); Petronikolou & Saridakis (2023) for examples of works exploring the implications of the Hubble tension for the nature of DE. The possibility of a more speculative “early dark energy” component, operative at very high redshift, is also a particularly interesting possibility in this context (see e.g. Poulin et al., 2019; Niedermann & Sloth, 2021; Braglia et al., 2020; Oikonomou, 2021a; Vagnozzi, 2021; Karwal et al., 2022; Benevento et al., 2022; Reeves et al., 2023; Poulin et al., 2023; Odintsov et al., 2023). Thus, probing the expansion history of the Universe at intermediate redshifts $z\approx 6-10$ constitutes a key opportunity to study the validity of $\Lambda$CDM model in an independent manner, and in an epoch which is otherwise extremely difficult to access (see also Gómez-Valent et al., 2023; Tutusaus et al., 2023, for recents work which highlighted the relevance of intermediate-redshift DE dynamics in the context of the Hubble tension).
@@
Among the various aspects of fundamental physics which can be put to test through measurements of the abundance of massive galaxies at very high redshift, of particular interest to the present work are the properties of DE. In particular, as explicitly demonstrated by one of us in Menci et al. (2020) using data from the CANDELS survey, the abundances of massive, high-redshift galaxies can provide key constraints on the DE equation of state (EoS) $w$. The reason is that $w$ controls the large-scale behaviour of DE at both background and perturbation level, thereby dramatically affecting the growth of density perturbations and the formation of cosmic structures. In this context, a number of recent works (see e.g. Lovell et al., 2022; Boylan-Kolchin, 2023; Forconi et al., 2023; Wang & Liu, 2022; Forconi et al., 2023) have demonstrated how initial JWST imaging data from NIRCam observations of the Cosmic Evolution Early Release Science (CEERS) program, which uncovered evidence of a surprisingly abundant population of very massive galaxies at extremely high redshifts $7\lesssim z\lesssim 10$, could severely challenge the $\Lambda$CDM scenario. Specializing to the implications for DE, it was argued by one of us in Menci et al. (2022) that these observations in fact exclude a significant portion of dynamical DE parameter space (including the point corresponding to $\Lambda$CDM), and favor models whose equation of state takes values $w<-1$ at some point after recombination, i.e. phantom models (Caldwell et al., 2003), see also Wang et al. (2023b) for further discussions.  [^5] For examples of other works showcasing the enormous potential of the early JWST observations in constraining various aspects of fundamental physics, see Biagetti et al. (2023); Haslbauer et al. (2022); Hütsi et al. (2023); Gandolfi et al. (2022); Maio & Viel (2023); Yuan et al. (2023); Dayal & Giri (2023); Ilie et al. (2023); Jiao et al. (2023); Parashari & Laha (2023); Hassan et al. (2023); Lei et al. (2023); Yoshiura et al. (2023); Padmanabhan & Loeb (2023); Lin et al. (2024); Su et al. (2023); Forconi et al. (2023); Gouttenoire et al. (2023); Guo et al. (2023); Huang et al. (2023); Bird et al. (2023); Wang et al. (2023b); Libanore et al. (2023); Wang et al. (2023a); van Putten (2024). Taking a step aside from new fundamental physics, a very important possibility of course is that the JWST results may call for a better understanding of galaxy formation, as discussed in a number of recent works (Ferrara et al., 2023; Qin et al., 2023; Pallottini & Ferrara, 2023; Wang et al., 2023; Pacucci et al., 2023). Although such results are still affected by uncertainties related to the assumed initial mass function (IMF) of the stellar populations (see e.g. Steinhardt et al., 2023a) and to the broadband selection criteria (especially for double-break sources, see for example Desprez et al., 2023), it is worth highlighting that recent spectroscopic observations in the redshift range $5\lesssim z\lesssim 6$ (Xiao et al., 2023) and $8\lesssim z\lesssim 10$ (Arrabal Haro et al., 2023) from the JWST First Reionization Epoch Spectroscopically Complete Observations (FRESCO) and CEERS surveys corroborate the above conclusions, while also placing them on a much stronger footing.
@@
The implications of the early JWST results for phantom DE are particularly interesting in view of the Hubble tension. It is now well understood that consistency with measurements of the acoustic angular scale at low redshifts from BAO requires that the sound horizon be lowered in the presence of a higher Hubble constant (Bernal et al., 2016; Addison et al., 2018; Lemos et al., 2019; Aylor et al., 2019; Schöneberg et al., 2019; Knox & Millea, 2020; Arendse et al., 2020; Efstathiou, 2021; Cai et al., 2022; Keeley & Shafieloo, 2023). This implies that phantom DE, an inherently late-time component which does not affect the sound horizon, cannot on its own completely solve the Hubble tension. Nevertheless, within the limits imposed by low-redshift data, phantom DE can at least partially alleviate the Hubble tension, in light of its ability to accommodate a higher Hubble constant while keeping the distance to the CMB, and thereby the acoustic angular scale observed in the CMB $\theta_{s}$, fixed (given that the sound horizon is unaffected, see Alestas et al., 2020, for a recent explicit discussion of parameter degeneracies in the presence of phantom DE). For this and other reasons, the possibility of phantom DE playing some role in the context of the Hubble tension is one which has been give serious consideration in the literature (see e.g. Zhao et al., 2017; Mörtsell & Dhawan, 2018; Li & Shafieloo, 2019; Vagnozzi, 2020; Alestas et al., 2021; Yang et al., 2021; Kumar, 2021; Teng et al., 2021; Heisenberg et al., 2023; Chudaykin et al., 2022; Sharma et al., 2023; Ballardini et al., 2023; Dahmani et al., 2023; Montani et al., 2024; da Costa et al., 2023, for recent discussions on the subject).
@@
Our discussions so far have been data-driven. However, theory considerations have a lot to offer to the discussion on viable DE models. Firstly, the tiny value of the positive CC $\Lambda$ required to explain observations is at severe odds with the value expected from theory considerations, when interpreted in terms of zero-point vacuum energy density of quantum fields: this is the well-known CC problem (Weinberg, 1989; Padmanabhan, 2003). Moving on to the simplest “quintessence” models for DE, based on a single, minimally coupled scalar field in the absence of higher derivative operators, and with canonical kinetic term (Wetterich, 1988; Peebles & Ratra, 1988; Ratra & Peebles, 1988; Wetterich, 1995; Caldwell et al., 1998), it is well known that these predict $w>-1$, and cannot therefore give rise to phantom DE, which instead requires a violation of the null energy condition (Vikman, 2005; Carroll et al., 2005; Nojiri et al., 2005; Oikonomou & Giannakoudi, 2022; Trivedi, 2023). It is also worth noting that, in the absence of additional ingredients, these “vanilla” quintessence models also worsen the Hubble tension, and one could therefore argue that they are observationally disfavored if the Hubble tension is to be taken seriously (Vagnozzi et al., 2018; Ó Colgáin et al., 2019; Ó Colgáin & Yavartanoo, 2019; Banerjee et al., 2021; Lee et al., 2022).
@@
Although most quintessence scalar field models feature a ground state with positive energy density, corresponding to a de Sitter (dS) vacuum, such a scenario has proven extremely difficult to construct in a controlled setting within string theory. In fact, it has been conjectured that string theory may be unable to accommodate dS vacua (Danielsson & Van Riet, 2018), as advocated by the swampland program (Vafa, 2005; Palti, 2019; Graña & Herráez, 2021), whose cosmological implications, particularly for inflation and DE, are far-reaching to say the least (see e.g. Obied et al., 2018; Agrawal et al., 2018; Achúcarro & Palma, 2019; Garg & Krishnan, 2019; Kehagias & Riotto, 2018; Kinney et al., 2019; Ooguri et al., 2019; Odintsov & Oikonomou, 2020; Oikonomou, 2021b; Cicoli et al., 2022) [^6] Two possible counterexamples to these conjectures are the KKLT (Kachru et al., 2003) and Large Volume Compactification (Balasubramanian et al., 2005) scenarios, although there is still no complete consensus on whether the resulting uplifted (meta)stable dS vacua are sufficiently long-lived to be able to agree with observations. On the other hand, anti-dS (AdS) vacua, which correspond to a negative CC (nCC), appear ubiquitously within string theory, and are among the best understood quantum gravity backgrounds by virtue of the AdS/CFT correspondence (Maldacena, 1998). Of course, a nCC $\Lambda<0$ with energy density $\rho_{\Lambda}<0$ is unable, on its own, to give rise to cosmic acceleration. However, an evolving DE component with positive energy density $\rho_{x}>0$ on top of a nCC can be consistent with the observed late-time acceleration, provided $\rho_{x}+\rho_{\Lambda}$ is positive around the present time, and amounts to about 70% of the total energy budget. Such a scenario, which could be interpreted in terms of a quintessence field whose potential features a negative minimum (AdS vacuum), is of great theoretical interest in light of the previous string-driven considerations. In fact, a number of recent works have explored similar scenarios in light of standard CMB, BAO, and SNeIa observations (see e.g. Cardenas et al., 2003; Poulin et al., 2018; Dutta et al., 2020; Visinelli et al., 2019; Ruchika et al., 2023; Di Valentino et al., 2021a; Calderón et al., 2021; Sen et al., 2022; Malekjani et al., 2023; Adil et al., 2023), finding that these models perform equally well as $\Lambda$CDM, or are even potentially statistically preferred.  [^7] Another related very interesting possibility which has been explored in the literature involves an AdS phase around recombination, as in the case of so-called AdS-early dark energy models (Ye & Piao, 2020a, b; Ye et al., 2021; Jiang & Piao, 2021; Ye et al., 2023b; Jiang et al., 2024; Jiang & Piao, 2022; Jiang et al., 2023; Ye et al., 2023a; Peng & Piao, 2024). The strong theoretical motivation behind such models, as well as their consistency with standard cosmological probes, strongly motivates further studies thereof, and in particular a comparison against other available observations, potentially in different redshift ranges, where the deviations from $\Lambda$CDM can be substantially more pronounced.
@@
In Adil et al. (2023), it was shown by some of us that a DE sector featuring an evolving component on top of a nCC can, within regions of parameter space consistent with standard cosmological observations, drastically alter the growth of structure at high redshifts with respect to $\Lambda$CDM to the point of potentially restoring concordance with the (photometric) JWST CEERS observations. We stress the word “potentially” as this earlier analysis was merely an exploratory one: in fact, besides only having explored a few benchmark points in parameter space, a careful treatment of corrections required for an accurate comparison to the JWST observations was lacking, as well as a more complete assessment of the complementarity with standard cosmological probes (the last two points were, in fact, both explicitly mentioned as motivation for a more detailed follow-up work in the closing paragraph of Adil et al., 2023). Given the enormous promise shown by a model featuring an evolving DE component on top of a nCC, it is our goal in the present work to revisit the model and go beyond the exploratory analysis of Adil et al. (2023), by providing a full comparison of such a model against the JWST CEERS observations, while also including information from standard cosmological probes. To place our results on a much more solid footing from the observational point of view, we also consider spectroscopic data from the JWST FRESCO survey (Xiao et al., 2023). Overall, our work confirms and reinforces the extremely promising conclusions reached earlier in Adil et al. (2023), and provides further motivation for exploring dark sectors featuring components with negative energy densities, while also highlighting once more the enormous potential held by observations of the abundance of massive galaxies at very high redshift in testing fundamental physics.

### v2

@@ Abstract
Early observations with the James Webb Space Telescope (JWST) have revealed the existence of an unexpectedly large abundance of extremely massive galaxies at redshifts $z\gtrsim 5$: these are in tension with the predictions not only of the standard $\Lambda$CDM cosmology, but also with those of a wide class of dynamical dark energy (DE) models, and are generally in better agreement with models characterized by a phantom behaviour. Here we consider a model, inspired by string theory and the ubiquity of anti-de Sitter vacua therein, featuring an evolving DE component with positive energy density on top of a negative cosmological constant, argued in an earlier exploratory analysis to potentially be able to explain the JWST observations. We perform a robust comparison of this model against JWST data, considering both photometric observations from the CEERS program, and spectroscopic observations from the FRESCO survey. We show that the model is able to accommodate the JWST observations, with a consistency probability of up to $98\%$, even in the presence of an evolving component with a quintessence-like behaviour (easier to accommodate theoretically compared to phantom DE), while remaining consistent with standard low-redshift probes. Our results showcase the potential of measurements of high-redshift galaxy abundances in tests of fundamental physics, and their complementarity with standard cosmological probes.
@@ 1 Introduction
The concordance $\Lambda$CDM cosmological model constitutes at present one of the simplest frameworks to describe the evolution of our Universe, and owes its success to its ability to describe a wide range of cosmological and astrophysical observations. One of the cornerstones of current cosmology is dark energy (DE) [1, 2, 3], a component with negative pressure and positive energy density driving the late-time acceleration of the Universe, whose discovery in 1998 led to one of the biggest paradigm shifts in physics [4, 5]: within $\Lambda$CDM, DE takes the form of a positive cosmological constant (CC) $\Lambda$ [6]. Over the past two and a half decades, the nature of DE has to a large extent been investigated by means of three classes of cosmological observations: low-redshift ($z\lesssim 2$) probes based on Type Ia Supernovae (SNeIa) and Baryon Acoustic Oscillations (BAO), the latter observed in the clustering of various tracers of the large-scale structure, such as galaxies; and high-redshift ($z\sim 1100$) probes based on the Cosmic Microwave Background (CMB): see Ref. [7] for a recent review on observational probes of DE.  [^1] Nevertheless, it is worth emphasizing that the sensitivity of the CMB to (late-time) DE is indirectly mostly a low-redshift (geometrical) one, through the effect of DE on the distance to the CMB itself (see e.g. the recent discussion in Ref. [8]), whereas the sensitivity of the CMB to DE perturbations and the dynamics thereof is quite limited, except possibly in the case of models which deviate substantially from the positive CC $\Lambda$, which are however ruled out by observations. By virtue of the significant efforts devoted to the study of observations, underlying theory, and systematics associated to these probes, CMB, BAO, and SNeIa are generally referred to as standard cosmological probes and enjoy a (rightly deserved) privileged status for what concerns cosmological tests of the expansion of the Universe and of fundamental physics, with particular regard to the nature of DE. However, a series of growing cosmological tensions which appear to challenge the $\Lambda$CDM model [9], and the fact that some of these probes are close to being systematics-limited, are making it imperative to start looking for new independent probes, whose complementarity and synergy with the standard CMB+BAO+SNeIa can greatly enrich the landscape of methods to study the Universe: in this sense, a number of so-called “emerging” cosmological probes are rapidly gaining momentum, as beautifully summarized in the recent review of Ref. [10] [^2] Of course tests of DE are not limited to cosmological scales. Various tests of DE on astrophysical, astronomical, and even local scales, have been considered over the past decades, see Refs. [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] for examples in this sense.
@@
In this context, the long-awaited recent observations of massive, distant galaxies delivered by the James Webb Space Telescope (JWST) [27] have opened a remarkable window not only onto astrophysics and the process of galaxy formation, but also onto novel cosmological tests of the contents of our Universe and more generally fundamental physics [28, 29, 30]. This is possible because the abundance of galactic haloes of dark matter (DM) mass $M$ at redshifts $z\approx 5-10$ is predicted to be exponentially sensitive to the growth factor of perturbations $D(z)$, while also being a strongly decreasing function of redshift, and one which depends crucially on the (model-dependent) time-redshift relation [31] [^3] Distant galaxies are also useful as a test of cosmological models through their ages, which of course cannot exceed the age of the Universe: see for instance Refs. [32, 33, 34, 35, 36, 37, 38] for recent works exploring the important of stellar and galactic ages as a test of cosmological models. On the other hand, if we denote by $\Omega_{b}$ and $\Omega_{m}$ the cosmic density parameters for baryons and the total matter component respectively (and with $f_{b}\equiv\Omega_{b}/\Omega_{m}$), the stellar mass of a galaxy $M_{\star}$ cannot exceed the maximal allowed baryonic content of the host halo $f_{b}M$. This implies that comparing the theoretically predicted cumulative comoving stellar number density of massive ($M_{\star}\approx 10^{10}-10^{11}\,M_{\odot}$) galaxies, $n(>M_{\star},z)$, with the observed abundance of galaxies of given stellar mass at redshifts $z\approx 5-10$, can potentially provide strong constraints on the expansion history and growth of cosmological perturbations, and therefore also on the nature of DE, in a way which is completely independent of the complex baryonic physics involved in the formation of galaxies. Needless to say, such constraints would be extremely valuable in light of their high complementarity to state-of-the-art determinations from CMB, BAO, and SNeIa, both in terms of range of redshifts/cosmic times probed, as well as potential systematics involved.
As hinted to earlier, the possibility of testing the expansion history of the Universe at such intermediate redshifts is particularly valuable in view of the tension between local and high-redshift determinations of the Hubble constant $H_{0}$ (see e.g. Refs. [39, 40, 41, 42, 43, 44, 45, 46, 47, 48] for recent reviews). At a significance of $\gtrsim 5\sigma$, the Hubble tension poses what is probably one of the most serious observational challenges to $\Lambda$CDM, and can potentially lead us to completely rethink the nature of DE.  [^4] With no claims as to completeness, see Refs. [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75] for examples of works exploring the implications of the Hubble tension for the nature of DE. The possibility of a more speculative “early dark energy” component, operative at very high redshift, is also a particularly interesting possibility in this context (see e.g. Refs. [76, 77, 78, 79, 80, 81, 82, 83, 84, 85]). Thus, probing the expansion history of the Universe at intermediate redshifts $z\approx 6-10$ constitutes a key opportunity to study the validity of $\Lambda$CDM model in an independent manner, and in an epoch which is otherwise extremely difficult to access (see also Refs. [86, 87] for recents work which highlighted the relevance of intermediate-redshift DE dynamics in the context of the Hubble tension).
@@
Among the various aspects of fundamental physics which can be put to test through measurements of the abundance of massive galaxies at very high redshift, of particular interest to the present work are the properties of DE. In particular, as explicitly demonstrated by one of us in Ref. [88] using data from the CANDELS survey, the abundances of massive, high-redshift galaxies can provide key constraints on the DE equation of state (EoS) $w$. The reason is that $w$ controls the large-scale behaviour of DE at both background and perturbation level, thereby dramatically affecting the growth of density perturbations and the formation of cosmic structures. In this context, a number of recent works (see e.g. Refs. [89, 90, 91, 92, 93]) have demonstrated how initial JWST imaging data from NIRCam observations of the Cosmic Evolution Early Release Science (CEERS) program, which uncovered evidence of a surprisingly abundant population of very massive galaxies at extremely high redshifts $7\lesssim z\lesssim 10$, could severely challenge the $\Lambda$CDM scenario. Specializing to the implications for DE, it was argued by one of us in Ref. [94] that these observations in fact exclude a significant portion of dynamical DE parameter space (including the point corresponding to $\Lambda$CDM), and favor models whose equation of state takes values $w<-1$ at some point after recombination, i.e. phantom models [95], see also Ref. [96] for further discussions.  [^5] For examples of other works showcasing the enormous potential of the early JWST observations in constraining various aspects of fundamental physics, see Refs. [97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 91, 113, 114, 115, 116, 96, 117, 118, 119]. Taking a step aside from new fundamental physics, a very important possibility of course is that the JWST results may call for a better understanding of galaxy formation, as discussed in a number of recent works [120, 121, 122, 123, 124, 125]. Although such results are still affected by uncertainties related to the assumed initial mass function (IMF) of the stellar populations (see e.g. Ref. [126]) and to the broadband selection criteria (especially for double-break sources, see for example, see Ref. [127]), it is worth highlighting that recent spectroscopic observations in the redshift range $5\lesssim z\lesssim 6$ and $8\lesssim z\lesssim 10$ [29, 30] from the JWST First Reionization Epoch Spectroscopically Complete Observations (FRESCO) and CEERS surveys corroborate the above conclusions, while also placing them on a much stronger footing.
@@
The implications of the early JWST results for phantom DE are particularly interesting in view of the Hubble tension. It is now well understood that consistency with measurements of the acoustic angular scale at low redshifts from BAO requires that the sound horizon be lowered in the presence of a higher Hubble constant [128, 129, 130, 131, 132, 133, 134, 135, 136, 137]. This implies that phantom DE, an inherently late-time component which does not affect the sound horizon, cannot on its own completely solve the Hubble tension. Nevertheless, within the limits imposed by low-redshift data, phantom DE can at least partially alleviate the Hubble tension, in light of its ability to accommodate a higher Hubble constant while keeping the distance to the CMB, and thereby the acoustic angular scale observed in the CMB $\theta_{s}$, fixed (given that the sound horizon is unaffected, see Ref. [138] for a recent explicit discussion of parameter degeneracies in the presence of phantom DE). For this and other reasons, the possibility of phantom DE playing some role in the context of the Hubble tension is one which has been give serious consideration in the literature (see e.g. Refs. [139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153] for recent discussions on the subject).
@@
Our discussions so far have been data-driven. However, theory considerations have a lot to offer to the discussion on viable DE models. Firstly, the tiny value of the positive CC $\Lambda$ required to explain observations is at severe odds with the value expected from theory considerations, when interpreted in terms of zero-point vacuum energy density of quantum fields: this is the well-known CC problem [154, 155]. Moving on to the simplest “quintessence” models for DE, based on a single, minimally coupled scalar field in the absence of higher derivative operators, and with canonical kinetic term [156, 157, 158, 159, 160], it is well known that these predict $w>-1$, and cannot therefore give rise to phantom DE, which instead requires a violation of the null energy condition [161, 162, 163, 164, 165]. It is also worth noting that, in the absence of additional ingredients, these “vanilla” quintessence models also worsen the Hubble tension, and one could therefore argue that they are observationally disfavored if the Hubble tension is to be taken seriously [166, 167, 168, 169, 170].
@@
Although most quintessence scalar field models feature a ground state with positive energy density, corresponding to a de Sitter (dS) vacuum, such a scenario has proven extremely difficult to construct in a controlled setting within string theory. In fact, it has been conjectured that string theory may be unable to accommodate dS vacua [171], as advocated by the swampland program [172, 173, 174], whose cosmological implications, particularly for inflation and DE, are far-reaching to say the least (see e.g. Refs. [175, 176, 177, 178, 179, 180, 181, 182, 183, 184]).  [^6] Two possible counterexamples to these conjectures are the KKLT [185] and Large Volume Compactification [186] scenarios, although there is still no complete consensus on whether the resulting uplifted (meta)stable dS vacua are sufficiently long-lived to be able to agree with observations. On the other hand, anti-dS (AdS) vacua, which correspond to a negative CC (nCC), appear ubiquitously within string theory, and are among the best understood quantum gravity backgrounds by virtue of the AdS/CFT correspondence [187]. Of course, a nCC $\Lambda<0$ with energy density $\rho_{\Lambda}<0$ is unable, on its own, to give rise to cosmic acceleration. However, an evolving DE component with positive energy density $\rho_{x}>0$ on top of a nCC can be consistent with the observed late-time acceleration, provided $\rho_{x}+\rho_{\Lambda}$ is positive around the present time, and amounts to about 70% of the total energy budget. Such a scenario, which could be interpreted in terms of a quintessence field whose potential features a negative minimum (AdS vacuum), is of great theoretical interest in light of the previous string-driven considerations. In fact, a number of recent works have explored similar scenarios in light of standard CMB, BAO, and SNeIa observations (see e.g. Refs. [188, 189, 190, 191, 192, 193, 194, 195, 196, 197]), finding that these models perform equally well as $\Lambda$CDM, or are even potentially statistically preferred.  [^7] Another related very interesting possibility which has been explored in the literature involves an AdS phase around recombination, as in the case of so-called AdS-early dark energy models [198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208]. The strong theoretical motivation behind such models, as well as their consistency with standard cosmological probes, strongly motivates further studies thereof, and in particular a comparison against other available observations, potentially in different redshift ranges, where the deviations from $\Lambda$CDM can be substantially more pronounced.
@@
In Ref. [209], it was shown by some of us that a DE sector featuring an evolving component on top of a nCC can, within regions of parameter space consistent with standard cosmological observations, drastically alter the growth of structure at high redshifts with respect to $\Lambda$CDM to the point of potentially restoring concordance with the (photometric) JWST CEERS observations. We stress the word “potentially” as this earlier analysis was merely an exploratory one: in fact, besides only having explored a few benchmark points in parameter space, a careful treatment of corrections required for an accurate comparison to the JWST observations was lacking, as well as a more complete assessment of the complementarity with standard cosmological probes (the last two points were, in fact, both explicitly mentioned as motivation for a more detailed follow-up work in the closing paragraph of Ref. [209]). Given the enormous promise shown by a model featuring an evolving DE component on top of a nCC, it is our goal in the present work to revisit the model and go beyond the exploratory analysis of Ref. [209], by providing a full comparison of such a model against the JWST CEERS observations, while also including information from standard cosmological probes. To place our results on a much more solid footing from the observational point of view, we also consider spectroscopic data from the JWST FRESCO survey [29]. Overall, our work confirms and reinforces the extremely promising conclusions reached earlier in Ref. [209], and provides further motivation for exploring dark sectors featuring components with negative energy densities, while also highlighting once more the enormous potential held by observations of the abundance of massive galaxies at very high redshift in testing fundamental physics.